### v2

# Reconstructing Materials Tetrahedron: Challenges in Materials Information Extraction

Abstract
Discovery of new materials has a documented history of propelling human progress for centuries and more. The behaviour of a material is a function of its composition, structure, and properties, which further depend on its processing and testing conditions. Recent developments in deep learning and natural language processing have enabled information extraction at scale from published literature such as peer-reviewed publications, books, and patents. However, this information is spread in multiple formats, such as tables, text, and images, and with little or no uniformity in reporting style giving rise to several machine learning challenges. Here, we discuss, quantify, and document these challenges in automated information extraction (IE) from materials science literature towards the creation of a large materials science knowledge base. Specifically, we focus on IE from text and tables and outline several challenges with examples. We hope the present work inspires researchers to address the challenges in a coherent fashion, providing a fillip to IE towards a materials knowledge base.
1 Introduction
Understanding a material’s behaviour requires knowledge about its composition, properties, processing and testing protocols, and microstructure—represented as the materials science (MatSci) tetrahedron (see Fig. 1). These different aspects of a material are reported by researchers in peer-reviewed publications, patents, and other scientific documents. Recently, there have several attempts to exploit the advances in machine learning (ML) and artificial intelligence (AI) towards automated information extraction (IE) from literatureKononova et al. (2021); Zaki et al. (2023a); Ravinder et al. (2021); Zaki et al. (2024). These include the development of materials specific language models Gupta et al. (2022); Huang and Cole (2022); Trewartha et al. (2022); Gupta et al. (2023), rule-based systems Swain and Cole (2016); Mavracic et al. (2021); Zaki et al. (2022b); Venugopal et al. (2021); Zaki et al. (2022c), IE from tables Gupta et al. (2023); Zhao and Cole (2022); Zhao et al. (2023), and IE from images Mukaddem et al. (2019); Zhang and Shao (2022); Zaki et al. (2023b, 2022a). The widely varying information expression styles in research papers makes the automated MatSci IE a challenging task. Most of the works have been focused on IE in a specific domain and hence the transferability to different materials is not explored. Moreover, there exists no consolidated work that explores the specific challenges associated with IE in MatSci and the gain associated with solving these challenges, which provides a clear direction to the researchers regarding the areas that require increased attention.
We thoroughly review MatSci articles to identify IE challenges towards completing the materials tetrahedron (see Fig. 1). We also highlight some of the major challenges toward the development of a “universal” MatSci knowledge base linking the extracted information from multiple sources and forms of data—structured, semi-structured, and unstructured. Indeed, there exist millions of scientific documents reporting information about various materials known to humans. Thus, the automated development of MatSci IE will lead to a rich knowledge base on materials. The outline of the paper is as follows: First, we explain the methodology of collecting papers for review and annotation process. Then, in the results and discussion sections, we investigate the proportion of each of the entities, such as composition, structure, properties, processing and testing conditions, reported in tables or text of the articles, followed by the challenges faced in their extraction. We quantify how frequently a challenge occurs to motivate researchers to gauge the amount of information that will be obtained after solving the respective challenges. We further identify the challenges in extracting and connecting the information from text and tables, and also among different tables belonging to the same MatSci research papers. Finally, we provide some guidelines for presenting machine and human-friendly tables that enable automated MatSci IE from research papers.
2 Methodology
To elucidate the challenges, we referred to a dataset of 2536 peer-reviewed publications on MatSci. This dataset is based on recent work on IE from tables Gupta et al. (2023), where the training data was annotated using distant supervision while the val and test data were annotated manually. Fig. 2 shows different sections of the paper where these different components are majorly reported. The statistics of each challenge were computed by randomly taking 50/100 tables from the manually annotated val and test dataset. In cases, where this was not applicable, we further performed manual annotation on additional 50 papers or 100 relevant tables, selected randomly from the corpus. For instance, for composition extraction, we randomly selected 100 composition tables from the manual annotation in the existing dataset. However, no such manual annotation was available for properties. For this problem, we selected 100 random property tables from the corpus and manually annotated the frequency of the challenges in property extraction. Note that all the challenges are based on manual annotation and hence are reliable than any ML-based technique such as distant supervision. Further, we manually analyzed tables or text for the occurrence of each of the entities such as composition, structure, and property. All the results and data associated with the annotation process are shared in the following link.
3 Results and Discussions
Figure 3 shows the percentage of papers reporting raw materials (precursors), compositions, properties, processing, and testing methods in text and tables. Note that the same information could be reported in both text and table and hence, the percentages may add to more than 100. Although 78% and 74% of papers had compositions in text and tables, respectively, an in-depth analysis revealed that only 33.21% of the total compositions were reported in the text, whereas 85.92% of compositions were present in tables. The overlap exists due to the same composition being mentioned in both text and tables. 82% articles report properties in tables (see Fig. 3). Processing and testing conditions are mostly reported in the text, while in 80% articles, precursors are mentioned in the text. In the following sections, we discuss these aspects in detail.
3.1 Composition extraction
Since the majority of the material compositions are reported in tables, we first discuss the challenges in extracting compositions from tables. This is followed by the discussion on IE from text.
3.1.1 Extracting compositions from tables
Here, we summarize the major challenges in IE from tables. To this extent, we investigated 100 randomly selected composition tables from the manually annotated data to report the frequency of occurrence of each challenge.
a. Variation in table structure and information content: An analysis of 100 random MatSci composition tables revealed that these tables do not follow any standard structure. Accordingly, following earlier schema proposed by Gupta et al. Gupta et al. (2023), composition tables can categorized into two broad categories—multi-cell composition (MCC) and single-cell composition (SCC) tables. These are further subdivided into tables containing complete information (CI) and partial information (PI). When the entire composition is written inside a single cell, it is classified as an SCC table, whereas when the composition is written across multiple cells of the table by reporting the value of each constituent (compounds or elements) of the composition in separate cells, it is defined to be an MCC table. If the table contains all the information regarding the constituents of the material, they are classified as CI tables (complete information). Alternatively, if only some of the constituents are mentioned in the table for the material, they are PI tables. In the latter case, we need to extend the analysis to the text of the article to extract the full composition. Fig. 4 illustrates all 4 types of tables Nazabal et al. (2003); Zaharescu et al. (2008); Narayanan and Shashikala (2015); Jestin et al. (2003). The most prevalent composition table types are MCC-CI (36%), followed by SCC-CI (30%). PI tables are less common, with 24% being present as MCC-PI and the remaining 10% as SCC-PI. Note that this distribution may vary significantly depending on the material types as well. For instance in alloys, it is a common practice to skip the major element while describing the composition in a table. In previous work by Gupta et al.. Gupta et al. (2023), while an F1 score of 78.21% and 65.41% have been achieve for extraction from SCC-CI and MCC-CI tables, respectively, an F1 score of only 51.66% has been achieved for extraction from MCC-PI. Extraction from SCC-PI is an open challenge.
b. Presence of nominal and experimental compositions: While the nominal composition is the amount of chemicals taken initially to prepare the material, analyzed/experimental composition refers to the actual composition of material obtained after analyzing the manufactured material (see Fig. 5(a)) Youngman et al. (2000); Petkov and Ewen (1999). Our analysis revealed that in 3% of the tables, both nominal and analyzed/experimental compositions are reported. These values are not reported in any fixed pattern, making it difficult to separate the nominal and analyzed compositions correctly after extraction.
c. Composition in table headers: Most tabular IE models like Tabbie Iida et al. (2021), DiSCoMaT Gupta et al. (2023) perform better when row/column headers contain appropriate information regarding its contents. In MatSci tables, the headers are mostly material IDs, compound names, properties, processing and testing labels, and the inner cells contain corresponding values. However, in 6% of the tables, we found that the compounds with their values were present in the heading, which makes it hard for the IE models to extract the desired information. For instance, SeGePb Jain et al. (2009) or = 10%, = 20%,… Minati et al. (2007) are column headers which contain both the compounds and corresponding concentration in the heading. 67% of these were SCC-CI, whereas the rest 33% were MCC-PI tables.
d. Compositions and related info inferred from other documents: In some tables, the details of the compositions studied are not explicitly mentioned; rather references to previous research publications which uses the same material is provided in the tables or their captions (see Fig. 5(b)). Thus, such a task requires the composition to be extracted from a previous work, which then should be combined with the property in the present work. Such references were seen in 11% of the tables Brasil and Meyer (2004); Chen et al. (2003).
e. Compounds inferred from material ID: 10% of the tables have IDs in which valuable information regarding the composition was present (see Fig. 6. For example, some of the material IDs have the molar ratios of compounds present in the material, while others have the composition written in an abbreviated form Nazabal et al. (2001). In such cases, compositions are not explicitly mentioned in the table. There are also tables where the composition of the materials is not specified; instead, their standard names are used as IDs. Such examples include Wollastonite or Diopside Schneider et al. (2000), which has a fixed chemical composition and therefore the ID can be used to infer the composition of the material.
f. Variables used to represent compounds: When a composition is expressed with variables such as TeO+15BO+15PO+LiO, where = 5, 10, 15, 20, 25 and 30 mol% Khafagy et al. (2008), it mostly denotes the variation of different compounds. However, in some articles, variables have been used to represent compound names instead of their values. One such example is REYAlCo where RE = Ce, Pr, Nd, Sm, Gd, Tb, Er, Sc Zhao et al. (2009). This scenario is observed in 1% of the tables.
3.1.2 Extracting compositions from text
Now, we report the challenges in extracting the compositions reported in the text of the MatSci research papers. We report our statistical findings based on each challenge’s occurrence frequency.
a. Different formats of compositions: The compositions in materials literature does not adhere to a predetermined pattern and encompasses various variations. This is in strict contrast to notations in chemistry where IUPAC nomenclature is used. Some notable examples are as follows.
1. "Erbium-doped glasses with the molar composition 40GeO.10SiO.25NbO.25KO, plus 0.1 to 4 mol% of ErO1.5, were prepared using mixtures of the respective oxides (99.99% purity), with exception of KO, which was added in the form of KCO" Santos et al. (2010).
2. "Bulk samples of (SeTe)Ag (0 x 4) system were prepared by conventional melt quenching technique. High-purity (99.999%) elements with appropriate atomic percentages were sealed in a quartz ampoule (length 100 mm and internal diameter 6 mm) in a vacuum of 10 - 5 mbar" Singh et al. (2012).
3. "The samples having chemical composition of 2(Ca,Sr,Ba)O–TiO–2SiO were examined. CaO, SrO and BaO contents in the samples were varied as shown in Table 1. RO% shows the molar percentage of CaO, SrO or BaO in total RO of CaO+SrO+BaO" Tsuzuku et al. (2002).
b. Extracting variable values in text: Extracting values from variables is challenging since the variable values are specified in different formats with some present only in the text. For instance, consider the following sentence from a peer-reviewed manuscript. Monteiro et al. (2013) A series of tellurite glasses with nominal composition (80-x)TeO–xGeO–10NbO–10KO, where x = 0, 10, 20, 30, 40, 50, 60, 70 and 80 mol%, hereafter named 8T0G, 7T1G, 6T2G, 5T3G, 4T4G, 3T5G, 2T6G, 1T7G and 0T8G, respectively, were doped with 0.2 to 4 mol% .
Here, the values representing the compositions and the respective variable names are present only in the text. A few instances of other formats of composition with variables are shown in Appendix 9.2.2(c).
c. Low recall in extracting compositions expressed with variables: 28% of the articles have compositions written with variables of which 28.57% does not provide any values for the variables in the text. Among the 71.53% where values are present, 40% of them do not mention the step size for the range of values taken by the variable. For example, consider the text representing a set of compositions as follows from a manuscript: x(0.75AgI:0.25AgCl):(1-x)(AgO:WO), where 0.1x1 in molar weight fraction. Agrawal et al. (2004) The step size of 0.1 is mentioned nowhere in the text but could be inferred from the composition table present in the paper. Therefore, extracting only from the text in such cases leads to more errors, and this can be resolved by connecting the variables to the correct composition table containing the variable.
d. Recognition of full forms and abbreviations: Instead of providing precise composition values, full forms are employed instead of abbreviations. Consider the following example.
"Lithium disilicate glass was prepared in 30 g quantity by heating stoichiometric homogeneous mixtures of lithium carbonate (99.0%), Synth, and silica (99.9999%), Santa Rosa, for 4 h at 1500 °C in a platinum crucible." Schröder et al. (2014). This text has the composition of the glass as lithium disilicate mentioned indirectly without clearly mentioning the percentages or numbers.
e. Unstable and irrelevant composition extraction: Unstable reagents and other irrelevant compositions which does not refer to the material are also identified as compositions due to a lack of robust parsers. is an unstable entity referring to the aluminium tetrahedral structure, while can be a composition. These undesired extractions can lead to a huge drop in the precision of the IE model, and separating them from the material composition is not an easy task. Only a domain expert with the help of the source article can confirm whether the extraction is relevant or not.
3.1.3 Extracting compositions from table and text jointly:
Extracting information from PI tables is more challenging than extracting from CI tables, as the incomplete information in the table regarding the composition should be inferred from the text. A detailed analysis of 50 PI tables revealed that 36% of the tables have unique challenges and are not ”regular”. To clarify this point further, we discuss some of these challenges below, while also defining a “regular” MCC-PI table in Fig 15.
a. Unusual variables used: Other than the common variables like , , , , and , we also encounter variables like , , and in 4% of the manuscripts. Distinguishing some of them such as S or Y is difficult as they are valid symbols for chemical elements as well Henry and Hill (2003).
b. Variables representing composition in text not found in tables: A generic way of extracting the composition when an arithmetic equation of the composition containing variables is mentioned in the text is to connect it to the variables located in the headings of the table and substitute it with values mentioned under it. Extraction becomes difficult if there is an absence or mismatch between the variable name in the table and the text, as shown in Fig. 7(a). We found 8% of the tables posing this challenge Yoshida et al. (2003).
c. Composition present across multiple columns: The composition of the material is spread across multiple columns/rows (instance depicted in Fig. 7(b) Kadono et al. (1999)), or the table does not follow any fixed orientation. This is observed among 4% of the PI tables.
d. Composition partly in the table and partly in text: Although PI tables contain the composition partly, it is expected that the complete information is available in the text. But in rare occurrences, as depicted by Fig. 7(c), we observe that only the remaining part of the composition, that isn’t mentioned in the table, is present in the text. This makes linking the parts of compositions in the text and tables challenging. Thus, extracting the whole composition is extremely difficult; a case seen in less than 1% of the PI tables.
e. Presence of multiple variables: 8% of the PI tables contains more than one variable, all of which need to be taken into account to extract the composition correctly. As discussed previously, variables can be of various forms, which might make the extraction of multiple variables a challenging task. Hayashi et al. (1998)
3.2 Extracting properties from tables
Until now, we focused on the extraction of compositions from tables. Now, we discuss the challenges with property extraction. To this extent, we analyzed 100 arbitrarily selected property tables. The observations based on this analysis are as follows.
a. Semantically similar row/column headers: 19% of the tables have similar abbreviations or headers with similar descriptions for different properties. For example, in Fig. 8(a), the headings of the columns are T, T, T, T, T, T Gu et al. (2002). Identifying the desired property by a predictor model or someone without domain knowledge can be difficult in this case.
b. Same property measured under different conditions: The same property can be measured with different techniques or under different conditions. Therefore, it is important to extract the correct contextual information related to the reported property. Some recurrent scenarios include witnessing tables with various refractive indexes (RIs) at different wavelengths Qiao et al. (2005) (see Fig. 11), glass transition temperatures at different heating rates Giridhar and Mahadevan (1998), or hardness at different testing loads. We encountered 9% of the property tables exhibiting this challenge.
c. Information in caption/footer instead of tables: Often, properties are mentioned with abbreviations in the headings of tables, which are semantically close to other properties (for example, Fig. 8(a)). The information regarding their abbreviation is commonly found in the caption or footer of the table. We observed 30% of the tables having this characteristic Scudino et al. (2005); Deparis et al. (2005). Further, 2% of the tables have no information on the properties units. However, these are found in the caption and footer of the tables Okuno et al. (2005). Hence, text from these sections might be handy for extracting our desired properties.
d. Property recorded under various acronyms: It is a common practice to record property names with their abbreviations. Some properties can have various abbreviations like density is represented with either or d, Young’s modulus with YM or E, and activation energy with E, Ae, or E.
e. Identical acronyms representing different entities: We encountered tables (see Fig. 8(b,c)) where the commonly used acronyms are used to represent different entities; not the usual property they generally represent. For example, ‘n’, which is mostly used to represent RI, is also used to represent equation parameters specific to the experiments. Another commonly seen instance is ‘d’ which is used to represent density Seshadri et al. (2014) and has also been used to represent fractal bond connectivity Abd El-Moneim (2014), lattice parameters and equation parameters. This suggests that using a string-matching IE algorithm can result in poor performance in such cases.
f. Range of values (min-max) given instead of mean value: In very few cases (< 1% tables), we encountered property values reported in range rather than a single value. For example, the values of T are reported in the range 930-945C Besson et al. (2000). Only a domain expert would know which value to take for a corresponding property between the min, max, or mean of the documented values. This might depend on the property or the application intended to be used, and will also be reflected in the IE algorithm.
3.3 Common for both composition and property extraction:
Thus far, we discussed the challenges faced during composition extraction in 3.1.1 and property extraction in 3.2 from tables. However, some challenges arise in either of these scenarios.
a. Same composition or property represented with different units: Occasionally, tables are presented where the essential information is recorded in multiple units within the same table in different columns/rows. Fig. 9 shows a composition table having composition in both mol% and wt% Konishi et al. (2000), and a property table having glass transition temperature (T) mentioned in both C and K Rocherullé et al. (1998). This can lead to duplication of the extracted data.
b. Multiple ways of reporting the same unit: Despite the well-known and accepted conventions for writing the SI units Thompson and Taylor (2008), research publications resort to multiple ways of reporting the same unit. For instance, for g/cm, several variations are observed in peer-reviewed publications such as gm/cm, g.cm, g/cm, gcm, g/cc, gm/cc, gw/cm, gm cc. Similar observations are made for kg/m, where variations such as kgm, kg/m, kg m are presented. Extracting the correct unit and normalizing it to a standard form is an essential task. Thus, while there are standard rules for writing SI units, it is observed that these are not followed strictly in scientific publications.
3.4 IE or manufacturing and characterizing materials
To identify the challenges in extracting precursors, processing and testing conditions, and material structure, we analyzed 50 arbitrarily selected papers from the dataset for reporting our findings.
a. Precursor extraction: A research paper generally investigates materials of a similar kind. Hence, it has to be assumed that all the materials are manufactured using the same precursors. In research papers where batch composition is mentioned in tables, the challenges are similar as mentioned in Section 3.1.1(b). In papers where researchers discuss the patented materials, they refer to them by their trademark name, for example, Pyrex, BOROSIL, Gorilla, etc. and hence their precursor information is not provided. However, papers discussing materials reported in previous publications, provide references to those papers reporting the required information in detail.
b. Processing conditions extraction: Processing conditions could be extremely non-linear and convoluted. Consider the set of sentences describing the processing conditions Roderick et al. (2001) as follows. “… powders were weighed and mixed thoroughly before being transferred to a 90 Pt/10 Rh crucible, heated at 320°C and maintained between 1000 and 1400°C depending on composition, for approximately 25 min. After annealing for approximately three hours, the glass was allowed to cool slowly to room temperature…”. Hence, the challenges here are to extract temperatures and duration for each process, like heating, annealing, and cooling, along with the environmental conditions and experimental apparatus. Sometimes, these conditions are also mentioned in the table (see Fig. 11), and their extraction poses similar challenges as described in Section 3.2(b).
c. Testing conditions extraction: The testing conditions mainly comprise the sample characteristics, dimensions, test name, instrument name, instrument settings, and testing variables like temperature, wavelength, load, etc. Consider the following excerpt from Sigoli et al. (2001): “The porous microstructure of the matrix was investigated by scanning electron microscopy (SEM) (JEOL JSM T330A), by infrared spectroscopy (IR) in a FT-IR spectrometer (Perkin Elmer Spectrum 2000), and by X-ray powder diffractometry (XRD) (Siemens D-5000). The phase separation process was investigated by Raman microscope. The room temperature Raman measurements were performed through Raman imaging microscope (Renishaw) system 3000, with the 632.8 nm He–Ne laser line for excitation”. The boldface text indicates the information to be extracted for obtaining a complete understanding of the testing process of a material. Fig. 11 lists different wavelengths at which a material is tested to obtain wavelengths. The challenges faced in IE for this case will be similar to the ones posed in Section 3.2(b).
Material structure: To study the structure of materials, researchers perform X-ray diffraction studies, obtain the Raman spectra, optical micrographs, and scanning electron micrographs depending upon the depth of detail about the material structure required. This information is mostly reported in figures and the figure description in the text provides some important details about the material’s structure. In the statement, “The Raman spectrum of the porous phase (Fig. 6(b)) shows only one band at 277 cm assigned to silica vibrations…”Sigoli et al. (2001), the information about Raman spectra is already shown in the graph, and the text mentions only critical findings.
To summarise, the extraction of precursors, processing, and testing conditions from text poses challenges related to named entity recognition and relation extraction, which requires the need for specialized datasets and model development. Further, the extracted entities should be linked with the respective materials. The challenges faced during IE from tables for these variables require overcoming similar challenges as explained earlier for composition (Section 3.1.1) and properties (Section 3.2)
3.5 MatSci Knowledge-base: Linking extracted information
The tetrahedron, as shown in Fig. 1 will be considered complete for a given material if its properties, processing, testing conditions, and raw materials required to manufacture are available. To this end, researchers need to link extracted compositions with these variables. These pose unique challenges as it requires linking information among different entities within the paper such as connecting different paragraphs of the paper, text with tables, or tables with other tables in the paper.
Material IDs are required to link information across multiple tables. For instance, in Fig. 12 Solvang et al. (2004), we obtain the composition of CAS1 from Table 1 and of this material from another table(Fig. 12 Table 2.). We detected 187 out of 2536 (7.37%) publications where inter-table IE is necessary and found difficulties in 81 of them while connecting the different components of the tetrahedron.
a. Different material IDs in different tables: The same materials have been reported with different IDs in different tables. 21 out of the 81 research papers (25.93%) have this challenge Romero et al. (2000); Martinelli et al. (2000).
b. Material IDs absent from tables: We detected a few MatSci papers where no IDs are present in the tables. There exist 23 out of 81 (28.40%) documents having this challenge Peitl and Zanotto (1999).
c. One of the tables does not contain material IDs: While connecting two tables, there are cases where IDs are mentioned only in one table Beggiora et al. (2003) (37 out of 81 (45.67%) papers with this challenge).
As we observe that material ID is a very important factor in connecting tables, we did an intensive analysis of the type of IDs that are reported in the tables (see Table. 1).
3.6 ID Analysis
As material ID is the key component in connecting materials from tables to text, across two different tables, or also across different sections of the text, we investigated arbitrarily selected 50 articles containing material IDs in the tables and recorded their semantic pattern to observe the semantics used by authors to refer to materials. We found that a majority of the authors prefer to use acronyms or self-made codes as IDs for referring to the materials, followed by natural numbers and standard material names, illustrated in Fig. 13. Material IDs are generally present at the beginning of the table and very rarely seen in the middle or end. Often, we come across tables having IDs which contain relevant information like the processing conditions of the material, or information about the state of the glass like amorphous or crystalline, or its composition, which are not separately mentioned in the table. These information are generally encoded as abbreviations, and extracting them can be challenging. In Table 1, we describe different cases along with the percentage occurrences. Note that the composition of the material should not be confused with its ID, as both are separate entities. An ID is expected to be shorter in length, most likely an acronym and unique to each material.
| Challenges in IE from material IDs | % of occurence |
| Composition info/doping conc. present only in IDs | 20 |
| ID Present in the middle | 2 |
| Multiple IDs Present for the same composition | 4 |
| State or structural info in ID | 2 |
| Info or references about the processing conditions | 8 |
| Same IDs but Different Composition | 4 |
| Article contains IDs Interconnected | 2 |
| Taken from other articles | 6 |
4 Guidelines for writing IE-friendly MatSci tables
Tables should be reported in such a way that automated extraction and the detection of the desired information are easy. Some of our suggestions are as follows (illustrated with Fig. 14, adapted from Kosuge et al. (1998)):
a. Use column orientation: Many IE algorithms which have been developed for tables have considered column orientation only. Also, we showed that 93% of the published tables are column-oriented. The following suggestions assume that we are following column orientation.
b. Use MCC-CI tables: Tables should have the components associated with a composition written in different cells. Moreover, the table should have the complete information of the material compositions (see Fig 4).
c. Use proper and descriptive headers: The headers should contain the chemical formula of the compounds or elements which make up the materials, along with the acronyms of the reported properties, with processing, and testing conditions. If precursors, processing, and testing conditions are common, they can be omitted from tables.
d. Use standard notations for units: Units should be mentioned in the column headers of the tables within brackets. Moreover, the standard notations for representing the SI unit should be consistently used.
e. All-in-one table: Prefer writing all the information of a particular material in a single table while following proper orientation. Following this will avoid the need for inter-table extraction.
f. IDs are mandatory: Material IDs are important to identify different materials mentioned in the tables and link them across tables and text. IDs should be mandatory for tables and written in the first column.
g. Consistent IDs: Material IDs should be formed as an acronym of its comprising constituents. They should be consistent in the whole article, i.e., there should not be more than one ID referring to the same material.
h. Table structure: Table should be of the structure [[Material ID], [C1], [C2], …, [P1], [P2], …]. ‘C’ denotes the constituting compound/elements that form the material. Sequence them so that their proportions are arranged in a descending manner. ‘P’ refers to the properties of the corresponding material.
i. Column/row-wise consistency: Each column or row should contain information only related to a particular entity mentioned in the heading or in the first row, respectively. Multiple tables or columns should not be concatenated into one.
j. Captions: All tables should have a clear, concise and descriptive caption. The table caption should clearly explain the acronyms used in the tables.
5 Conclusion and future work
The literature is replete with IE challenges and algorithms to extract information about materials. However, there exists no study which quantifies how much benefit can be obtained if a particular challenge is solved. In this paper, we have identified and quantified several unresolved challenges present in IE for every aspect of the MatSci tetrahedron. Specifically, we pointed out the locations in a MatSci research paper, where each piece of information on the MatSci tetrahedron of a given material is reported. Further, we outlined the challenges associated with IE and linking them to build the MatSci KB. We hope this extensive analysis will motivate researchers to focus on the challenges in the field, giving an insight into the gain associated with each of these challenges. This will also enable the researchers to identify the right problems to focus on based on the desired outcome. Finally, we provided recommendations towards an IE-friendly table format, which will enhance the automated extraction of the desired information and improve the tabular understanding of the researchers. Such concerted efforts are required to streamline the reporting in MatSci articles, thereby accelerating IE for materials discovery.
6 Author Contributions
Hira, K - Conceptualization, Methodology, Software, Validation, Investigation, Data Curation, Writing - Original Draft.
Zaki, M - Conceptualization, Methodology, Validation, Investigation, Visualization, Data Curation, Writing - Original Draft.
Sheth, D. - Investigation, Data Curation, Writing - Original Draft.
Mausam - Conceptualization, Writing - Review & Editing, Supervision, Project administration.
Krishnan, N.M.A. - Conceptualization, Writing - Review & Editing, Supervision, Project administration.
7 Conflicts of interest
There are no conflicts to declare.
8 Acknowledgements
N. M. Anoop Krishnan acknowledges the funding support received from BRNS YSRA (53/20/01/2021-BRNS), Google, and Intel Labs. Mohd Zaki acknowledges the funding received from the PMRF award by the Government of India. Mausam acknowledges grants by Google, IBM, Verisk, and a Jai Gupta chair fellowship. He also acknowledges travel support from Google and Yardi School of AI travel grants. The authors thank the High Performance Computing (HPC) facility at IIT Delhi for computational and storage resources.